### v1

# Gradient Diffusion: A Perturbation-Resilient Gradient Leakage Attack

Abstract
Recent years have witnessed the vulnerability of Federated Learning (FL) against gradient leakage attacks, where the private training data can be recovered from the exchanged gradients, making gradient protection a critical issue for the FL training process. Existing solutions often resort to perturbation-based mechanisms, such as differential privacy, where each participating client injects a specific amount of noise into local gradients before aggregating to the server, and the global distribution variation finally conceals the gradient privacy. However, perturbation is not always the panacea for gradient protection since the robustness heavily relies on the injected noise. This intuition raises an interesting question: is it possible to deactivate existing protection mechanisms by removing the perturbation inside the gradients? In this paper, we present the answer: yes and propose the Perturbation-resilient Gradient Leakage Attack (PGLA), the first attempt to recover the perturbed gradients, without additional access to the original model structure or third-party data. Specifically, we leverage the inherent diffusion property of gradient perturbation protection and construct a novel diffusion-based denoising model to implement PGLA. Our insight is that capturing the disturbance level of perturbation during the diffusion reverse process can release the gradient denoising capability, which promotes the diffusion model to generate approximate gradients as the original clean version through adaptive sampling steps. Extensive experiments demonstrate that PGLA effectively recovers the protected gradients and exposes the FL training process to the threat of gradient leakage, achieving the best quality in gradient denoising and data recovery compared to existing models. We hope to arouse public attention on PGLA and its defense.
Xuan Liu* HKPolyU Siqi Cai* WHUT Qihua Zhou SZU
Song Guo HKUST Ruibin Li HKPolyU Kaiwei Lin WHUT
1 Introduction
With the latest success in real-life applications of various data-based generative models, the privacy and information security of these models are under increasing scrutiny[2]. Federated Learning (FL), as a newly proposed privacy preservation distributed machine learning paradigm, collaboratively trains a global model on the parameter server by aggregating computed updates from local clients [18], is widely used in various training scenarios[19, 21, 27], for example, Gboard and Medical Diagnosis System, where data security and privacy are highly significant[30]. Specifically, FL protects users’ privacy by sharing only the model parameters or gradients but not the raw data [15]. Existing FL protocol designs have been demonstrated to be vulnerable to adversaries both inside and outside the system, jeopardizing data privacy and system robustness [17]. Among these, gradient leakage attacks have shown serious privacy hazards [10, 16, 35]. To better protect the private information of the clients from malicious attackers during the training and inference periods, protection of shared gradients becomes an almost essential step [23]. Among existing gradient protection methods, gradient perturbation, as an extensively used privacy-preserving method, has already been frequently applied in FL scenarios efficiently [20] (schematic diagram shown in Fig. 1).
Figure 1: Threat model. The FL training process is threatened by gradient leakage attacks, where the attacker can intercept the exchanged gradients $\nabla W$ to recover the private training data. Previous work often protects the gradients by injecting perturbation into the gradients to form $\nabla W^{\prime}_{N}$ and $\nabla W^{Protect}$. Our PGLA removes the perturbation injected in the protected gradients via the diffusion-based denoising model.
Considering that gradient data is the result of linear transformations applied to the original data, such as image data, we posit that if the original data is stable and predictable, then the corresponding gradient data will also be stable and predictable. Moreover, we discovered that the essence of gradient perturbation protection is akin to the diffusion process applied to inherently structurally stable gradient data. Based on these two points, we believe that compared to other models, diffusion models [12] have a natural applicability for gradient perturbation denoising. In this paper, we reveal the natural diffusion properties of gradient perturbation protection and propose PGLA: Perturbation-resilient Gradient Leakage Attack, the first general gradient diffusion attack that is capable of rendering various kinds of perturbations on gradients (e.g. differential privacy [26], certain layer representation perturbation [25], dynamic perturbation [29]) nearly invalid. Based on the structure of obtaining protected shared gradients, our approach involves constructing a surrogate downstream task model for the target attack model. Without access to the original model structure or third-party datasets, we utilize the surrogate gradient data supply model to provide training data for our Gradient Diffusion Model. The trained Gradient Diffusion Model enables us to approximate the original shared gradients of the target clients from the perturbed privacy-preserving gradients obtained by malicious attackers. PGLA embeds the perturbation protection mechanism into the reverse diffusion process and regulates privacy-preserving gradients denoising node as well as the diffusion forward and reverse time steps with the perturbation level $M$ as an adaptable parameter. In summary, the key contributions of our work are:
  • We disclose the natural diffusion process in general gradient perturbation mechanisms by introducing perturbation adaptive parameter $M$.
  • We propose PGLA: Perturbation-resilient Gradient Leakage Attack, which successfully breaks the bottleneck that existing gradient leakage attacks cannot effectively leak privacy under gradient perturbation protection.
  • We present and demonstrate the vulnerability of gradient perturbation protection under diffusion denoising and the suitability of the diffusion model on gradient morphology data.
  • We design the first practical gradient diffusion attack strategy (including surrogate model construction, adjusting gradients to build diffusion training dataset, etc.), and conduct experiments under general perturbation protection FL system. Results show that PGLA achieves the best gradients denoising quality and privacy leakage ability on commonly used image datasets.
2 Preliminary
Federated Learning (FL) provides comparative privacy guarantees for model training which does not require the original data and the participating clients are free to enter or leave [17]. Considering a general FL system with homogeneous architectures. Regarding the gradient protection issue in the FL training process, gradient perturbation is one of the most widely used mechanisms. Here, we present the key preliminary definitions of gradient perturbation and gradient attacks that are pertinent to our work.
Table 1: Explanation of Necessary Notations.
$\nabla W$ Original gradients without any perturbation
$\nabla W^{\prime}$ FL shared gradients protected by perturbation
$\nabla W^{R}$ Recovered gradients of PGLA
$\nabla W_{j}$ Surrogate gradients for PGLA training
$(x_{j},y_{j})$ $Random_{(data,label)}$ for gradients extract
$N$ Number of clients in FL
$F_{i}$ Client’s model (i: $1\rightarrow N$)
$W^{s}$ Surrogate model weight parameter
$F^{s}$ Surrogate model
$X_{0}$ Data input of PGLA training
$\alpha_{t},\beta_{t}$ PGLA forward process variances
$f_{\theta}$ PGLA noise prediction function
$M$ Noise perturbation scale on gradients
$\varepsilon$ DP privacy budget
$\nabla s$ DP sensitivity ($\nabla s_{c}-client$,$\nabla s_{s}-sever$)
$\sigma_{dpc/s}$ DP noise scale (Standard Deviation)
2.1 Differential Privacy (DP)
Generally, DP is defined based on the concept of adjacent database and has been applied in various practical areas in Artificial Intelligence to protect privacy information through adding specific perturbation, e.g. Google’s RAPPORT [8] and large-scale graph data publishing [4]. Tracing back to the definition of Differential Privacy, in this paper we discuss both $\varepsilon-DP$ and $(\varepsilon,\delta)-DP$ [5, 6, 1] as our sample attack background. Here, we present two core definitions first.
Definition 1. A randomized mechanism $M$: $D\rightarrow R$ with domain $D$ and range $R$ satisfies $\varepsilon$ differential privacy if for any two adjacent inputs d, d’ $\in D$ and for any subset of outputs $S\subseteq R$, and it holds that:
$Pr[M(d)\in S]\leq e^{\varepsilon}Pr[M(d^{\prime})\in S].$ (1)
Definition 2. A randomized mechanism $M$: $D\rightarrow R$ with domain $D$ and range $R$ satisfies $(\varepsilon,\delta)$ differential privacy if for any two adjacent inputs d, d’ $\in D$ and for any subset of outputs $S\subseteq R$, and it holds that:
$Pr[M(d)\in S]\leq e^{\varepsilon}Pr[M(d^{\prime})\in S]+\delta.$ (2)
$\delta$-approximation is preferably smaller than $1/|d|$. Note that we usually apply Laplace perturbation for the definition in Eq. (1). However, as to the definition in Eq. (2), the noise perturbation mechanism needs to be Gaussian Mechanisms, which can adapt both $\varepsilon$ and $\delta$. According to Theorem A.1 in Gaussian Mechanisms [7] to ensure the given Gaussian noise distribution $N(0,\sigma^{2}_{dp})$ well preserves $(\varepsilon,\delta)-DP$, the noise scale should satisfy :
$\sigma_{dp}$$\geq$ $c_{dp}\nabla s/\varepsilon,\quad\varepsilon\in(0,1),$ (3)
$c_{dp}$$\geq$ $\sqrt{2ln(1.25/\delta)},$ (4)
where $\nabla s=max_{D,D^{\prime}}||S(D)-S(D^{\prime})||$ and $s$ represents the real-value function. $c_{dp}$ can be considered as a hyperparameter used to define the DP boundary.
2.2 FL with Differential Privacy (FL-DP)
FL-DP protects the shared model parameters or gradients between clients and the server during the FL training process by applying Local Differential Privacy (LDP [33, 26]) and (or) Differential Privacy Stochastic Gradient Descent algorithm (DPSGD [1, 34]).
Taking the local participates site for example, the main actions for FL-DP can be divided into four steps: setting DP noise mechanism, local gradient clipping, local gradient perturbation, and uploading the protected parameters to the server [28]. FL involves multiple participants, making composition theorem in DP necessary when applying noise. Among the composition DP methods available, including Simple Composition, Advanced Composition [7], and Moments Accountant [1], we utilize Simple Composition for the following discussion. In the case where $M_{i}$ satisfies $(\varepsilon,\delta)-DP$, the composition ($M_{1},M_{2},\cdots,M_{k}$) satisfies ($\sum_{i=1}^{k}\varepsilon_{i},\sum_{i=1}^{k}\delta_{i}$). Aligning with the state-of-art FL-DP framework [28], We applied Simple Composition with the following noise calculation formulas for local client gradient perturbation, considering both $\varepsilon-DP$ (using the Laplace mechanism) and $(\varepsilon,\delta)-DP$ (using the Gaussian mechanism):
$\textit{Laplace}:\sigma_{dpc}$ $=$ $\nabla s_{c}\times 1/\varepsilon,$ (5)
$\textit{Gaussian}:\sigma_{dpc}$ $=$ $\nabla s_{c}\times\sqrt{2ln(1.25/\delta)}/\varepsilon,$ (6)
where $\nabla s_{c}$ represents the sensitivity and can be formulated as $\nabla s_{c}=\frac{2C}{m}$. Note that $C$ is the clipping threshold for bounding $||\nabla W_{i}||\leq C$, where $\nabla W_{i}$ denotes the unperturbed gradient from the $i-th$ client and $m$ is the minimum size of the local datasets.
From the parameter server site, referring to noising before model aggregation FL (NbAFL) [28], the standard deviation $\sigma_{dps}$ of Gaussian noises that are added to $\nabla W_{server}$ (aggregated gradients at the server to be broadcast to clients) in an $N-client$ FL system is as follows:
$\sigma_{dps}=\left\{\begin{array}[]{cl}\frac{c_{dp}\nabla s_{s}}{\varepsilon}&\quad when\quad T_{dps}>L_{dpc}\sqrt{N}\\ 0&\quad when\quad T_{dps}\leq L_{dpc}\sqrt{N}\\ \end{array}\right.$ (7)
$\nabla s_{s}=\frac{2C}{m}\times\frac{\sqrt{T_{dps}^{2}-L_{dpc}^{2}N}}{N}.$ (8)
The variables $T_{dps}$ and $L_{dpc}$ represent the aggregation times and exposures of uploaded parameters from each client under the FL-DP system, respectively.
2.3 FL with Perturbation Protection (FL-PP)
In addition to DP, FL-PP includes a range of alternative gradient perturbation techniques, which do not need to satisfy DP concepts on similarity and can employ noise to effectively obscure private data. (e.g. adding random global noise to gradients, adding noise to specific layers, and dynamic noise addition) [32].
2.4 Gradient Inversion Attack (GradInv)
GradInv is a prevalent method in gradient leakage attacks and aims to steal the client’s privacy information in the FL system. The primary idea for the majority of GradInv to recover original information is minimizing the distance between the dummy gradient $\nabla W_{\xi}$ and the original gradient $\nabla W$ while updating the random data $x_{\xi}$ and label $y_{\xi}$ until the optimized results $\nabla W^{*}_{\xi}$ and $(x^{*}_{\xi},y^{*}_{\xi})$ are close enough to the original ones. The key formulation can be described as:
$\quad min||\nabla W_{\xi}-\nabla W||:(x_{\xi},y_{\xi})\rightarrow(x^{*}_{\xi},y^{*}_{\xi}).$ (9)
Previous works utilize Peak Signal-to-Noise Ratio (PSNR) of the recovered images to evaluate the performance of GradInv, with the threshold for the success of such attacks typically hinging on the degree of detail discernible to the human eye in the reconstructed private images[22, 9, 35, 16]. To effectively safeguard privacy information against GradInv, a plethora of gradient perturbation methods have been extensively employed [35].
[…]
3 Methodology
PGLA is the first Diffusion Attack Method focusing on the gradient data structure that can be applied to multiple kinds of gradient perturbation protection in FL through adaptive parameters setting on both forward and reverse process of the Gradient Diffusion Model employed in PGLA. The threat models targeted by PGLA encompass various forms of gradient perturbation. Meanwhile, the extremely similar noise mechanism on gradient perturbation protection and diffusion Markovian process provides a mathematical necessity for PGLA to realize an efficient privacy attack on gradient perturbation protection. Overall, our PGLA method (shown in Fig. 2) can be summarized in the following four steps:
(1) Get Protected Gradients. Honest-but-curious malicious attackers steal the Shared Perturbed Gradients $\nabla W^{\prime}$ during the FL training process by hiding on the server side or waiting on the way of parameter sharing (Fig.1).
Algorithm 1 Gradients Extracted for PGLA Training
1:  Stolen Protected Gradients: $F_{i}\rightarrow\nabla W^{\prime}$;
2:  Construct surrogate Model: $\nabla W^{\prime}\rightarrow F^{s}$;
3:  $j=0$; $\epsilon\sim N(\textbf{0},\textbf{I})$;
4:  for $iteration=1$ to $Length_{randomdataset}$ do
5:     $\nabla W_{j}=\partial l(F^{s}(x_{j};W^{s}),y_{j})/\partial W^{s}$;
6:     Save $\nabla W_{j}$;
7:     $j=j+1$;
8:  end for
[…]
Before delving into Step (3) and Step (4), two configurations that recur in the subsequent steps are defined in advance:
[Gradient Adjustment]: Gradients each with total size $L$ ($L=L_{\nabla W_{Layer1}}+L_{\nabla W_{Layer2}}+...+L_{\nabla W_{LayerN}}$) are adjusted into $1\times g\times g$ ($g=L+P$: $g=$ minimum integer satisfies $g^{2}>L$; $P=$ 0-padding size) before feeding into Gradient Diffusion Model for training or inference to adapt gradient diffusion process. The adjustment is only related to Diffusion procedures, gradients are restored to their original structure and size before Gradient-Based Attacks and Evaluations.
[M-Adaptive Process]: Adaptive parameter $M$ is inserted into $\nabla W^{\prime}$ before inference to ensure the starting time step is appropriately positioned to maximize the denoising capability of Gradient Diffusion Model (Eq. (16) $\sim$ Eq. (18)).
[…]
Algorithm 2 Gradient Diffusion Model Training
1:  if With Condition $\nabla W^{\prime}$ then
2:     $X_{0}=(\nabla W^{\prime},\nabla W_{j})$;
3:  else
4:     $X_{0}=\nabla W_{j}$;
5:  end if
6:  Repeat:
7:   $X_{0}\sim p(X_{0})$; $t\sim Uniform(1\rightarrow T)$; $\epsilon\sim N(\textbf{0},\textbf{I})$;
8:   Take a gradient descent step on: $\nabla_{\theta}||\epsilon-f_{\theta}(\sqrt{\gamma_{t}}X_{0}+\sqrt{1-\gamma_{t}}\epsilon,t)||^{2}$;
9:  until converged;
[…]
Figure 3: Visualization of Markovian gradient diffusion process. $M$ is the noise scale of perturbation, which is set as the adaptive parameter in [M-Adaptive Process]. PGLA Train and PGLA Inference correspond to Algorithm 2 and Algorithm 3 respectively.
PGLA’s Gradient Diffusion Model is inspired by DDPM[12], to be specific:
In Step (3), refer to Algorithm 2 lines 4 to 7, we set $t\in(1,T)$ as the time steps of Gaussian noise addition. $T$ is the total time step of the forward Markovian diffusion process. $\alpha_{t}(0<\alpha_{t}<1)$ denote the adaptive variables at each iteration and $\gamma_{t}=\prod_{t=0}^{t}\alpha_{t}$. With the above settings, the forward process ($q$: with no learnable parameters) of PGLA’s Gradients Diffusion Model is:
$q(X_{1}|X_{0})$ $=$ $N(X_{1};\sqrt{\alpha_{1}}X_{0},(1-\alpha_{1})I),$ (10)
$q(X_{t}|X_{0})$ $=$ $N(X_{t};\sqrt{\gamma_{t}}X_{0},(1-\gamma_{t})I).$ (11)
Given $(X_{0},X_{t})$, $X_{t-1}$ can be modeled as:
$q(X_{t-1}|X_{0},X_{t})=N(X_{t};\mu_{t},\sigma_{t}I).$ (12)
Algorithm 3 Generate Original Gradient
1:  if Known Noise Scale $M$ then
2:     $\textbf{c}=\frac{1}{\sqrt{1+M^{2}}}$;
3:  else
4:     $\quad\textbf{c}\in(0,1)$;
5:  end if
6:  $X_{T^{\prime}}\sim\textbf{c}\nabla W^{\prime}$;
7:  for $t=T^{\prime},...,1$ do
8:     if $t>1$ then
9:        $\textbf{z}\sim N(\textbf{0},\textbf{I})$;
10:     else
11:        $\quad\textbf{z}=\textbf{0}$;
12:     end if
13:     $X_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(X_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\gamma_{t}}}f_{\theta}(X_{t},t))+\sqrt{1-\alpha_{t}}\textbf{z}$;
14:  end for
15:  Return $X_{0}$;
16:  $\nabla W^{R}\leftarrow X_{0}$; (For Further Gradient-Based Attack)
If insert the stolen Shared Perturbed Gradients $\nabla W^{\prime}$ or constructed Surrogate Perturbed Gradients $\nabla W^{s}_{perturbed}$ as training condition then input $X_{0}=(condition,\nabla W_{j})$.If not to train with the condition, then input $X_{0}=\nabla W_{j}$.Through algebraic calculation, $\mu_{t}$ and $\sigma_{t}$ in Eq. (12) can be simplified for further usage in the reverse training process[24] as:
$\mu_{t}$ $=$ $\frac{\sqrt{\gamma_{t-1}}(1-\alpha_{t})}{1-\gamma_{t}}X_{0}+\frac{\alpha_{t}(1-\gamma_{t-1})}{1-\gamma_{t}}X_{t},$ (13)
$\sigma^{2}_{t}$ $=$ $\frac{1-\gamma_{t-1}}{1-\gamma_{t}}(1-\alpha_{t}).$ (14)
In the reverse process $(p)$ of PGLA’s Gradient Diffusion training, the objective function $f_{\theta}$ which is trained to predict noise vector $\epsilon$ is modeled as:
$E_{X_{0},\epsilon}[\frac{(1-\alpha_{t})^{2}}{2\sigma^{2}_{t}\alpha_{t}(1-\gamma_{t})}||f_{\theta}(\sqrt{\gamma_{t}}X_{0}+\sqrt{1-\gamma_{t}}\epsilon,t)-\epsilon||^{2}].$ (15)
In Step (4), refer to Algorithm 3 lines 1 to 4, two different original gradient generation processes are chosen depending on whether or not attackers know the Noise Scale of Perturbation of the Threat Model. Take FL with Gaussian Differential Privacy (Eq. (6) in Preliminary) as an example, consider an experienced and clever attacker who may know the DP Privacy Budget $\varepsilon$ and the probability of information leakage $\delta$ ($\delta$ is likely to be set as $10^{-5}$ from usual practice in DP), combining with the sensitivity $\nabla S$ (which can be estimated if the attacker has some previous information with the target model training dataset), the noise scale can be calculated or estimated to an approximate value $M$ ($M$ defined as the adaptive parameter in [M-Adaptive Process]). So, $\nabla W^{\prime}$ can be modeled as:
$\nabla W^{\prime}=\nabla W+MN(0,I).$ (16)
Considering the forward Markovian gradient diffusion process in Step 3 Eq. (10) &\&& Eq. (11), the relation between stolen perturbed gradients $\nabla W^{\prime}$ and original gradients $\nabla W$ can be remodeled as:
$\frac{1}{\sqrt{1+M^{2}}}\nabla W^{\prime}=\frac{1}{\sqrt{1+M^{2}}}\nabla W+\frac{M}{\sqrt{1+M^{2}}}N(0,I).$ (17)
Recall that our target is to generate $\nabla W^{R}$, so $\frac{1}{\sqrt{1+M^{2}}}\nabla W^{\prime}$ should be considered as $X_{t}$ in the inverse process of Gradient Diffusion Model, while $\nabla W$ stands for $X_{0}$. Correspondingly, during the construction of the recovered gradient $\nabla W^{R}$, the inverse time steps can be set as $T^{\prime}$. The relationship between $T^{\prime}$ and $M$ is:
$\frac{1}{1+M^{2}}=\prod_{t=0}^{T^{\prime}}\alpha_{t}.$ (18)
Since $\gamma_{t}=\prod_{t=0}^{t}\alpha_{t}$ have been predefined and calculated during the forward Markovian diffusion process, $T^{\prime}$ can be fixed in an approximate range $T^{\prime}\in(T^{\prime}_{-},T^{\prime}_{+})$, where $\prod_{t=0}^{T^{\prime}_{-}}\alpha_{t}<\frac{1}{1+M^{2}}<\prod_{t=0}^{T^{\prime}_{+}}\alpha_{t}$ and $T^{\prime}_{-or+}$ are positive integers less than the total forward noise addition time step $T$. On the other hand, if $M$ can not be estimated, which means the attacker knows nothing about the noise scale. Since $\frac{1}{1+M^{2}}\in(0,1)$, the input value of the inference process can simply be set as $c\nabla W^{\prime}$ where $c\in(0,1)$ to adapt the Markovian forward process.
Also, if conditions allow, $c$ can be modeled and predicted by a separate Machine Learning model according to the specific requirement of attackers.
4 Experiments
We assess the performance of the PGLA within the FL-PP framework, which encompasses the FL-DP approach.
4.1 Experimental Setups.
Variant Attack Models Three PGLA variant models: PGLA (trained with only unperturbed surrogate gradients), Condition PGLA (trained with both perturbed gradients and unperturbed surrogate gradients), and Non-Adaptive PGLA (without [M-Adaptive Process]: no perturbation scale $M$ as adaptive parameter during gradient diffusion process) are presented during experiments.
Datasets In the subsequent experimental analyses, we employ the datasets MNIST, CIFAR100, and STL10 as client privacy datasets, which also serve as the ground truth for privacy leakage evaluation. We extract the unperturbed original gradients ($\nabla W$) of the aforementioned three datasets from the local training model of the target client as the reference benchmark of gradient denoising under the FL-PP paradigm.
Evaluation and Boundary In the context of evaluating Gradient Denoising (Sec. 4.2 $\sim$ 4.3), we employ the cosine similarity $CosSimilar_{g}$ and the Average Peak Signal-to-Noise Ratio $PSNR_{g}$ as metrics to assess the quality of the Recovered Gradients $\nabla W^{R}$ compared to the Original Gradients $\nabla W$. Higher values of the two metrics indicate better accuracy in the original gradient recovery.
To evaluate the Privacy Leakage Capability (Sec. 4.4), we utilize the Image Average Peak Signal-to-Noise Ratio $PSNR_{i}$ and the Label Recovered Accuracy $LRA$. These metrics are employed to assess the fidelity of the recovered images and the accuracy of the recovered labels, respectively, to the original images and ground truth labels. The boundary of Privacy Leakage Attack is aligned with previous works (Sec. 2.4): the attack is considered successful if human visual perception can discern the requisite information from the recovered images.
4.2 Gradients Denoising under FL-DP
Among FL-DP, we choose the NbAFL framework [28] (Noising before model aggregation FL, details refer to Sec. 2.2) as the threat model in this experiment due to its widespread adoption. To ensure a comprehensive evaluation of the effectiveness of the PGLA, we train the three PGLA variant models, as well as non-diffusion denoising models such as NBNet [3], SS-BSN [11], and AP-BSN [14], using gradients extracted from the surrogate downstream task model trained on the FashionMnist. Following that, we utilize the trained models to denoise the shared perturbed gradients intercepted by the attacker from the target clients. These target clients have locally trained downstream task models using privacy datasets (MNIST, CIFAR100, and STL10). The gradients are protected using the NbAFL[28] before being shared with the server.
Figure 4: The visualization of gradients denoising on the MNIST dataset.
Based on the experimental results presented in Table 2, it is evident that PGLA showcases superior denoising ability for perturbed gradients under FL-DP, with an average Cosine Similarity exceeding 0.992 and a PSNR of 37.68 on the three datasets. This represents a significant improvement of over $27\%$ compared to non-diffusion methods. Fig.4 visualizes the denoising process of gradients with different methods, we can also observe that PGLA has a better denoising effect compared to NBNET, while SS-BSN and AP-BSN destroy the original structure of the gradient due to their over-denoise effect.
Further analysis reveals that among the three PGLA variant models, the performance can be ranked as follows: PGLA(Condition)$>$ PGLA $>$ PGLA (Non-Adaptive). This ranking demonstrates both M-Adaptive Process and conditional training enhance the denoising and generation performance of PGLA’s Gradient Diffusion Model. It suggests that selecting an appropriate PGLA variant model based on available prior information can improve attack performance.
Table 2: Overall results on gradients denoising under FL-DP. Threat model setting: DP-($\varepsilon$, $\delta$) = (2,$10^{-5}$), with Gaussian perturbation; NA: Non-Adaptive; C: Condition.
Model MNIST CIRAF100 STL10
$CosSimilar_{g}$ $PSNR_{g}$ $CosSimilar_{g}$ $PSNR_{g}$ $CosSimilar_{g}$ $PSNR_{g}$
Noised Gradient 0.992 35.74 0.979 27.51 0.979 27.23
NBNet[3] 0.792 18.07 0.764 15.48 0.768 15.61
SS-BSN[11] 0.971 29.82 0.845 22.85 0.845 22.84
AP-BSN[14] 0.968 29.61 0.892 24.01 0.893 24.12
PGLA 0.996 38.76 0.990 30.42 0.990 30.01
PGLA(NA) 0.995 38.59 0.990 30.39 0.990 29.87
PGLA(C) 0.996 38.78 0.990 30.53 0.993 30.22
4.3 Gradients Denoising under FL-PP
To evaluate the generalization capability of PGLA on gradient denoising, we constructed two different types of noise (Laplace and Gaussian) randomly applied to each layer of the original gradients within the FL-PP framework. The training and inference processes of PGLA variant models, and non-diffusion denoising models (NBNet [3], SS-BSN [11], and AP-BSN [14]) are the same as the above experiments on gradients denoising under FL-DP. In contrast to attacks specifically tailored for the FL-DP framework, the gradient denoising experiments on FL-PP showcase PGLA’s ability to effectively handle various types of perturbations and adapt to different magnitudes of perturbation.
Referring to the experimental results illustrated in Fig. 5, it is observed that when subjected to Laplace perturbation, PGLA variant models exhibit an average improvement of $21.6\%$ in PSNR and $9.2\%$ in Cosine Similarity. Similarly, under Gaussian perturbation, PGLA achieves an average enhancement of $21.3\%$ in PSNR and $9.1\%$ in Cosine Similarity. These findings provide compelling evidence for the robustness and stability of the PGLA and its variant models.
Figure 5: Gradients denoising under FL-PP: Gaussian and Laplace perturbed gradients denoising performance of $PSNR_{g}$ and $CosSimilar_{g}$ via different perturbation magnitudes. The same noise magnitude parameter $\varepsilon$ as in FL-DP is used to represent different levels of perturbation. Perturbation decreases, $\varepsilon$ increases. (Private dataset: STL10)
4.4 Privacy Leakage Capability
The comparison of overall privacy leakage capability from perturbed gradients of PGLA and typical Gradient Leakage Attacks (GRNN [22], IG [9], and DLG [35]) under FL-DP, compares Image Average Peak Signal Noise Ratio $PSNR_{i}$ and Label Recovered Accuracy $LRA$ of the recovered local clients’ privacy information. Local clients’ private training datasets are MNIST, CIFAR100, and STL10.
Table 3: Privacy leakage capability of PGLA variant models and traditional Gradients Leakage Attacks under FL-DP ($\delta$ = $10^{-5}$, $\varepsilon$ = 1, 5, 10). NA: Non-Adaptive; C: Condition; Gray marker: PGLA outperforms the highest result of traditional ones.
Datasets Model $\bm{\varepsilon}$=1 $\bm{\varepsilon}$=5 $\bm{\varepsilon}$=10
$PSNR_{i}$ LRA $PSNR_{i}$ LRA $PSNR_{i}$ LRA
MINIST GRNN[22] 14.70 1.000 16.54 1.000 20.19 1.000
IG[9] 3.721 - 5.164 - 6.067 -
DLG[35] 0.222 0.676 8.025 0.752 17.08 0.909
PGLA 17.87 0.851 24.17 0.895 33.09 0.927
PGLA(NA) 17.39 0.851 24.14 0.889 32.98 0.925
PGLA(C) 20.87 0.887 28.39 0.924 35.14 0.957
CIFAR100 GRNN[22] 18.24 1.000 20.61 1.000 21.15 1.000
IG[9] 3.684 - 5.073 - 5.971 -
DLG[35] 0.112 0.769 7.673 0.833 18.57 0.896
PGLA 18.25 0.883 19.04 0.902 21.32 0.902
PGLA(NA) 18.15 0.871 17.98 0.884 21.06 0.891
PGLA(C) 18.69 0.901 21.02 0.902 23.49 0.908
STL10 GRNN[22] 12.24 1.000 16.24 1.000 20.98 1.000
IG[9] 4.323 - 6.082 - 7.851 -
DLG[35] 0.038 0.783 5.673 0.815 17.25 0.851
PGLA 12.36 0.806 16.30 0.857 19.77 0.889
PGLA(NA) 12.28 0.801 15.19 0.849 19.76 0.887
PGLA(C) 12.98 0.816 20.30 0.872 22.77 0.906
Figure 6: Visualization results on data recovery. Comparisons on ground truth images and corresponding recovered privacy images from PGLA variant models and commonly used traditional gradient leakage attacks. ($\delta$ = $10^{-5}$; $\varepsilon=10$)
According to the numerical experimental results shown in Table 3 and the visualization results shown in Fig. 6, PGLA variant models exhibit a substantial superiority over traditional gradient leakage attacks in terms of private image leakage. On average, there is an approximately $209\%$ increase in the recovered image $PSNR_{i}$ when using PGLA. It’s worth noting that GRNN can always accurately recover the label, keeping its LRA metric at 1.000. The reason is that GRNN has uniquely designed a robust generative network specifically to recover labels, while other methods obtain the label by updating the image and label jointly with a single model during the backward phase, leading to joint image and label recovery. Moreover, upon examining the private image recovery procedures of PGLA variant models and traditional gradient leakage attacks illustrated in Fig. 7, it becomes evident that the attacks incorporating PGLA not only achieve considerably improved accuracy in the ultimate reconstruction of private images compared to conventional approaches but also showcase notable advantages in terms of attack iteration rounds and speed (PGLA (condition) $>$ PGLA $>$ PGLA (Non-Adaptive)). Above all, PGLA variant models demonstrate significant advantages in recovering private images.
Figure 7: Comparison of the private image recovery procedures to the iterations between PGLA variant models and traditional Gradient Leakage Attack methods (DLG[35]).
4.5 Attack Duration
To comprehensively evaluate the attack efficiency of PGLA, we compared the average inference time of different denoising models under FL-PP. The training and inference procedures of the variant models of PGLA, as well as the non-diffusion denoising models, are configured to be consistent as in Sec. 4.3.
The experimental results presented in Table 4 indicate that PGLA variant models exhibit a relative advantage in terms of gradient denoising speed, surpassing non-diffusion methods by an average improvement of $32.7\%$ in inference time.
Limitations: The experiments demonstrate that PGLA can reconstruct original gradients from noise-perturbed gradients without requiring knowledge of the type and magnitude of perturbation noise, as well as FL clients’ local model. However, PGLA is not effective in attacking perturbations that are not based on gradient diffusion (noise perturbation) such as representation perturbation. In addition, the overall effectiveness of PGLA in reconstructing privacy datasets is also bounded by the selected subsequent Gradient Leakage Attacks.
Table 4: Gradient denoising average inference time of PGLA variant models and non-diffusion denoising models under FL-PP. (Private dataset: STL10, $\varepsilon$ = 0.5 $\sim$ 10)
Model NBNet [3] SS-BSN [11] AP-BSN[14] PGLA PGLA(NA) PGLA(C)
Times (s) 2.653 26.15 7.138 6.834 6.148 6.917
Defense Strategies: Possible defense strategies against PGLA can be approached from preserving the privacy of original data, preserving the target attack models, and shared gradients protection. PGLA effectively leakages privacy data by attacking the perturbed shared gradients, which suggests the possible defense approaches against PGLA should either be non-perturbation gradient protection methods or non-gradient privacy-preserving methods. For instance, processing of the participant’s local training model in depth and layers to weaken the performance of Gradient Leakage Attacks, encryption on original local privacy data before training, and various cryptography-based protections on shared gradients [13, 32, 23]. Cryptography-based approaches always cause more computation time and storage [31], thus lightweight defense strategies that can serve as substitutes for invalid perturbation protection still need future explorations.
5 Conclusion
This paper makes the first attempt to investigate the diffusion property of the widely-used perturbation-based gradient protection in Federated Learning (FL). To reveal the potential vulnerability, we propose a novel Perturbation-Resilient Gradient Leakage Attack (PGLA), an effective attack paradigm to deactivate the perturbation protection methodology by leveraging the denoising capability of diffusion models. Based on PGLA, we wish to enhance public consciousness on the issues of perturbation denoising and gradient leakage in FL applications.