### v2

# Mjölnir: Breaking the Shield of Perturbation-Protected Gradients via Adaptive Diffusion

Abstract
Perturbation-based mechanisms, such as differential privacy, mitigate gradient leakage attacks by introducing noise into the gradients, thereby preventing attackers from reconstructing clients’ private data from the leaked gradients. However, can gradient perturbation protection mechanisms truly defend against all gradient leakage attacks? In this paper, we present the first attempt to break the shield of gradient perturbation protection in Federated Learning for the extraction of private information. We focus on common noise distributions, specifically Gaussian and Laplace, and apply our approach to DNN and CNN models. We introduce Mjölnir, a perturbation-resilient gradient leakage attack that is capable of removing perturbations from gradients without requiring additional access to the original model structure or external data. Specifically, we leverage the inherent diffusion properties of gradient perturbation protection to develop a novel diffusion-based gradient denoising model for Mjölnir. By constructing a surrogate client model that captures the structure of perturbed gradients, we obtain crucial gradient data for training the diffusion model. We further utilize the insight that monitoring disturbance levels during the reverse diffusion process can enhance gradient denoising capabilities, allowing Mjölnir to generate gradients that closely approximate the original, unperturbed versions through adaptive sampling steps. Extensive experiments demonstrate that Mjölnir effectively recovers the protected gradients and exposes the Federated Learning process to the threat of gradient leakage, achieving superior performance in gradient denoising and private data recovery.
Introduction
Figure 1: Threat model. The FL training process is threatened by gradient leakage attacks, where the attacker can intercept the exchanged gradients $\nabla W$ to recover the private training data. Previous work often protects the gradients by injecting perturbation into the gradients to form $\nabla W^{\prime}_{N}$ and $\nabla W^{Protect}$. Our Mjölnir removes the perturbation injected in the protected gradients via the adaptive diffusion process.
Federated Learning (FL) is a distributed machine learning paradigm that facilitates collaborative model training without directly transmitting raw training data. Instead, it aggregates gradients or parameters shared among clients to build a global model (McMahan et al. 2017; Gong et al. 2024; Wu et al. 2023; Zhang et al. 2024). This approach preserves the privacy of raw data by keeping it within its originating domain, addressing concerns associated with traditional centralized data processing. However, FL is vulnerable to gradient inversion attacks (Geng et al. 2023; Liu et al. 2023; Zhu, Liu, and Han 2019), in which adversaries can potentially reconstruct sensitive user data from the shared gradients. This vulnerability has spurred significant research into gradient protection techniques (Tan et al. 2024a; Rodríguez-Barroso et al. 2023). Gradient Perturbation, such as differential privacy (DP), injecting noise into gradients to enhance privacy, has been proven to be an effective strategy for safeguarding data in FL scenarios (Wei et al. 2020; Ouadrhiri and Abdelhadi 2022; Wang, Hugh, and Li 2024).
Perturbation-based gradient protection achieves its goal by adding noise to the gradients. If a method is developed to effectively eliminate this noise, the protective mechanism of this approach becomes ineffective. Diffusion Model (Ho, Jain, and Abbeel 2020; Gong 2023) has the natural applicability for denoising perturbation, and is a potential approach to attack perturbation-based gradient protection. We hold the above idea based on two points: 1) Considering that gradient is the result of linear transformations applied to the training data, such as image data, we posit that if the original data is stable and predictable, then the corresponding gradient data will also be stable and predictable. 2) The essence of gradient perturbation protection is akin to the diffusion process applied to inherently structurally stable gradient data.
In this paper, we present the first attempt to break the shield of gradient perturbation protection in FL based on the diffusion model. We reveal the natural diffusion properties of gradient perturbation protection and propose Mjölnir [^1] Mjölnir is the hammer wielded by Thor and is renowned for its incredible power to break any shield., a perturbation-resilient gradient leakage attack method and the first general gradient diffusion attack (schematic diagram shown in Fig. 2) that is capable of rendering various kinds of perturbations on gradients (e.g. Differential Privacy (Truex et al. 2020), certain layer representation perturbation (Sun et al. 2021), dynamic perturbation (Wei et al. 2021)) nearly invalid. As the first attempt, we focus on common noise distributions, specifically Gaussian and Laplace, and apply our approach to DNNs and CNNs. Our method involves constructing a surrogate model for the target attack model to obtain protected shared gradients. In the absence of the original model structure and third-party datasets, we use the surrogate gradient data supply model to generate training data for our gradient diffusion model. This trained gradient diffusion model allows us to approximate the original shared gradients from the perturbed, privacy-preserving gradients obtained by attackers. Mjölnir integrates the perturbation protection mechanism into the reverse diffusion process, adjusting the privacy-preserving gradient denoising nodes and regulating the diffusion’s forward and reverse time steps with the perturbation level $M$ as an adaptable parameter.
In summary, our key contributions include:
  • We disclose the natural diffusion process in general gradient perturbation mechanisms and introduce a perturbation adaptive parameter $M$ to adaptively adjust the diffusion step size according to the degree of perturbation.
  • We propose Mjölnir, the first practical gradient diffusion attack strategy to recover the perturbed gradients, without additional access to the original model structure and third-party data, which breaks the bottleneck that existing gradient leakage attacks cannot effectively leak privacy under gradient perturbation protection.
  • We demonstrate the vulnerability of gradient perturbation protection under the Mjölnir adaptive diffusion denoising process. Experimental results under the general perturbation protection FL system show that Mjölnir achieves the best gradient denoising quality and privacy leakage ability on commonly used image datasets.
Background and Related Work
FL with Perturbation Protection (FL-PP)
FL-PP includes a variety of alternative gradient perturbation techniques, such as adding random global noise to gradients, noise to specific layers, and dynamic noise addition (Zhang et al. 2022). Among these, FL with Differential Privacy (FL-DP) is the most widely used method (Shi et al. 2022; Chen et al. 2023; Tan et al. 2024b).
Generally, DP is defined based on the concept of adjacent database and has been applied in various practical areas in Artificial Intelligence to protect privacy information through adding specific perturbation, e.g. Google’s RAPPORT (Erlingsson, Pihur, and Korolova 2014) and large-scale graph data publishing (Ding et al. 2021). FL-DP protects the shared model parameters or gradients between clients and the server during the FL training process by applying Local Differential Privacy (LDP (Zhao et al. 2021; Truex et al. 2020)) and (or) Differential Privacy Stochastic Gradient Descent algorithm (DPSGD (Abadi et al. 2016; Zhou et al. 2023)). In this paper we discuss both $\varepsilon-DP$ and $(\varepsilon,\delta)-DP$ (Dwork et al. 2006a, b; Abadi et al. 2016) as our sample attack background (Shi et al. 2022; Chen et al. 2023; Tan et al. 2024b):
Assumption 1. A randomized mechanism $M$: $D\rightarrow R$ with domain $D$ and range $R$ satisfies $\varepsilon$ differential privacy if for any two adjacent inputs d, d’ $\in D$ and for any subset of outputs $S\subseteq R$, and it holds that:
$Pr[M(d)\in S]\leq e^{\varepsilon}Pr[M(d^{\prime})\in S]$ (1)
Assumption 2. A randomized mechanism $M$: $D\rightarrow R$ with domain $D$ and range $R$ satisfies $(\varepsilon,\delta)$ differential privacy if for any two adjacent inputs d, d’ $\in D$ and for any subset of outputs $S\subseteq R$, and it holds that:
$Pr[M(d)\in S]\leq e^{\varepsilon}Pr[M(d^{\prime})\in S]+\delta$ (2)
$\delta$-approximation is preferably smaller than $1/|d|$. We usually apply Laplace perturbation for the definition in Eq. (1). However, as to the definition in Eq. (2), the noise perturbation mechanism needs to be Gaussian Mechanisms, which can adapt both $\varepsilon$ and $\delta$. Refer to Theorem A.1 in Gaussian Mechanisms (Dwork and Roth 2014), with $\nabla s=max_{D,D^{\prime}}||S(D)-S(D^{\prime})||$ ($s$ is the real-value function) and $c_{dp}$ denotes the hyperparameter used to define the DP boundary, to ensure Gaussian noise distribution $N(0,\sigma^{2}_{dp})$ well preserves $(\varepsilon,\delta)-DP$, the noise scale should satisfy:
$\sigma_{dp}$$\geq$ $c_{dp}\nabla s/\varepsilon,\quad\varepsilon\in(0,1)$ (3)
$c_{dp}$$\geq$ $\sqrt{2ln(1.25/\delta)}$ (4)
Taking the local participates site for example, the main actions for FL-DP can be divided into four steps: setting DP noise mechanism, local gradient clipping, local gradient perturbation, and uploading the protected parameters to the server (Wei et al. 2020). FL involves multiple participants, making composition theorem in DP necessary when applying noise. Among the composition DP methods available, including Simple Composition, Advanced Composition (Dwork and Roth 2014), and Moments Accountant (Abadi et al. 2016), we utilize Simple Composition for the following discussion. In the case where $M_{i}$ satisfies $(\varepsilon,\delta)-DP$, the composition ($M_{1},M_{2},\cdots,M_{k}$) satisfies ($\sum_{i=1}^{k}\varepsilon_{i},\sum_{i=1}^{k}\delta_{i}$). Aligning with the state-of-art FL-DP framework (Wei et al. 2020; Dwork et al. 2006a, b; Abadi et al. 2016; Shi et al. 2022; Chen et al. 2023; Tan et al. 2024b), We applied the following assumptions for noise calculation in local client gradient perturbation, considering both $\varepsilon$-DP (using the Laplace mechanism) and $(\varepsilon,\delta)$-DP (using the Gaussian mechanism):
$\textit{Laplace}:\sigma_{dpc}$ $=$ $\nabla s_{c}\times 1/\varepsilon$ (5)
$\textit{Gaussian}:\sigma_{dpc}$ $=$ $\nabla s_{c}\times\sqrt{2ln(1.25/\delta)}/\varepsilon$ (6)
where $\nabla s_{c}$ is the sensitivity and can be formulated as $\nabla s_{c}=\frac{2C}{m}$. $C$ is the clipping threshold for bounding $||\nabla W_{i}||\leq C$, where $\nabla W_{i}$ denotes the unperturbed gradient from the $i-th$ client and $m$ denotes the minimum size of local datasets.
Gradient Inversion Attack (GradInv)
GradInv is a prevalent method in gradient leakage attacks and aims to steal the client’s privacy information in the FL system. The primary idea for the majority of GradInv to recover original information is minimizing the distance between the dummy gradient $\nabla W_{\xi}$ and the original gradient $\nabla W$ while updating the random data $x_{\xi}$ and label $y_{\xi}$ until the optimized results $\nabla W^{*}_{\xi}$ and $(x^{*}_{\xi},y^{*}_{\xi})$ are close enough to the original ones. The key formulation can be described as:
$\quad min||\nabla W_{\xi}-\nabla W||:(x_{\xi},y_{\xi})\rightarrow(x^{*}_{\xi},y^{*}_{\xi})$ (7)
Previous works utilize Peak Signal-to-Noise Ratio (PSNR) of the recovered images to evaluate the performance of GradInv, with the threshold for the success of such attacks typically hinging on the degree of detail discernible to the human eye in the reconstructed private images (Geiping et al. 2020; Zhu, Liu, and Han 2019; Liu et al. 2023).
Methodology
Figure 2: Mjölnir Overview. After intercepting the exchange gradients protected by unknown perturbation ($\nabla W^{\prime}$) from clients, the attacker will (1) leverage the invariance of gradient data structure before and after perturbation to construct the surrogate model; (2) feed random image dataset into the surrogate model to extract surrogate clean gradients; (3) flatten the surrogate gradients and pad them to the appropriate size $g^{2}=P+L$ (g is the minimum integer satisfies $g^{2}>L$) to create a surrogate gradient ($\nabla W^{s}$) dataset for training Gradient Diffusion Model (if additional conditions are chosen to guide the training process, joint ($\nabla W^{s}$,$\nabla W^{s}_{perturbed}$) or ($\nabla W^{s}$,$\nabla W^{\prime}$) as the training dataset, where $\nabla W^{s}_{perturbed}$ denote surrogate gradients applied with known perturbation; if no conditions are needed, directly use $\nabla W^{s}$); (4) use the trained Gradient Diffusion Model to denoise $\nabla W^{\prime}$ to generate the recovered gradient $\nabla W^{R}$. From $\nabla W^{R}$, the attacker can recover clients’ privacy information.
Mjölnir is the first Diffusion Attack Method focusing on the gradient data structure that can be applied to multiple kinds of gradient perturbation protection in FL through adaptive parameters setting on both forward and reverse process of the Gradient Diffusion Model employed in Mjölnir. The threat models targeted by Mjölnir encompass various forms of gradient perturbation. Meanwhile, the extremely similar noise mechanism on gradient perturbation protection and diffusion Markovian process provides a mathematical necessity for Mjölnir to realize an efficient privacy attack on gradient perturbation protection. Overall, our Mjölnir method (shown in Fig. 2) can be summarized in four steps:
Step (1) Get Protected Gradients. Honest-but-curious malicious attackers steal the Shared Perturbed Gradients $\nabla W^{\prime}$ during the FL training process by hiding on the server side or waiting on the way of parameter sharing (Fig.1).
Algorithm 1 Gradients Extracted for Mjölnir Training
1:Stolen Protected Gradients: $F_{i}\rightarrow\nabla W^{\prime}$;
2:Construct surrogate Model: $\nabla W^{\prime}\rightarrow F^{s}$;
3:$j=0$; $\epsilon\sim N(\textbf{0},\textbf{I})$;
4:for $iteration=1$ to $Length_{randomdataset}$ do
5:     $\nabla W_{j}=\partial l(F^{s}(x_{j};W^{s}),y_{j})/\partial W^{s}$;
6:     Save $\nabla W_{j}$;
7:     $j=j+1$;
Step (2) Construct Surrogate Model. Construct Surrogate Gradients Data Supply Model $F^{s}$ from the data structure of Shared Perturbed Gradients $\nabla W^{\prime}$ stolen by the attacker that can output the same gradient data structure as the target attacked client’s local model (Algorithm 1, Fig. 2).
Before delving into Step (3) and Step (4), two configurations that recur in the subsequent steps are defined:
[Gradient Adjustment]: Gradients each with total size $L$ ($L=L_{\nabla W_{Layer1}}+L_{\nabla W_{Layer2}}+...+L_{\nabla W_{LayerN}}$) are adjusted into $1\times g\times g$ ($g=L+P$: $g=$ minimum integer satisfies $g^{2}>L$; $P=$ 0-padding size) before feeding into Gradient Diffusion Model for training or inference to adapt gradient diffusion process. The adjustment is only related to Diffusion procedures, gradients are restored to their original structure and size before further Gradient-Based Attacks and Evaluations.
[M-Adaptive Process]: Adaptive parameter $M$ is inserted into $\nabla W^{\prime}$ before inference to ensure the starting time step is appropriately positioned to maximize the denoising capability of Gradient Diffusion Model (Eq. (14) $\sim$ Eq. (16)).
Algorithm 2 Gradient Diffusion Model Training
1:if With Condition $\nabla W^{\prime}$ then $X_{0}=(\nabla W^{\prime},\nabla W_{j})$;
[…]
3:Repeat:
4:$X_{0}\sim p(X_{0})$; $t\sim Uniform(1\rightarrow T)$; $\epsilon\sim N(\textbf{0},\textbf{I})$;
5: Take a gradient descent step on: $\nabla_{\theta}||\epsilon-f_{\theta}(\sqrt{\gamma_{t}}X_{0}+\sqrt{1-\gamma_{t}}\epsilon,t)||^{2}$;
6:until converged;
Step (3) Train Gradients Diffusion Model. Construct our Gradient Diffusion Model that takes into account the level of knowledge about the attacked model (e.g., whether the level of perturbation noise or the type of noise is known). Conduct [Gradient Adjustment] to Clean Gradients $\nabla W^{s}$ extracted from Surrogate Model $F^{s}$ to build a training gradient dataset to train our Gradient Diffusion Model (Fig. 2, Fig.3). Algorithm 2 demonstrates a detailed training process.
Step (4) Recover Original Gradients. The stolen Shared Perturbed Gradients $\nabla W^{\prime}$ are put into our trained Mjölnir Gradient Diffusion Model after [Gradient Adjustment] and [M-Adaptive Process] to generate the Recovered Gradients $\nabla W^{R}$ for further Gradient-Based Attack to get certain privacy information based on various demands (Algorithm 3, Fig. 2 and Fig. 3).
Figure 3: Visualization of Markovian gradient diffusion process. $M$ is the noise scale of perturbation, which is set as the adaptive parameter in [M-Adaptive Process]. Mjölnir Train and Mjölnir Inference correspond to Algorithm 2 and Algorithm 3 respectively.
Algorithm 3 Generate Original Gradient
1:if Known Noise Scale $M$ then $\textbf{c}=\frac{1}{\sqrt{1+M^{2}}}$;
[…]
3:$X_{T^{\prime}}\sim\textbf{c}\nabla W^{\prime}$;
4:for $t=T^{\prime},...,1$ do
5:     if $t>1$ then $\textbf{z}\sim N(\textbf{0},\textbf{I})$;
[…]
7:     $X_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(X_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\gamma_{t}}}f_{\theta}(X_{t},t))+\sqrt{1-\alpha_{t}}\textbf{z}$;
8:return $X_{0}$; $\nabla W^{R}\leftarrow X_{0}$: $\triangleright$ For Further GradInv;
Mjölnir’s Gradient Diffusion Model is inspired by DDPM(Ho, Jain, and Abbeel 2020), to be specific:
In Step (3) Train Gradients Diffusion Model, refer to Algorithm 2 lines 4 to 6, we set $t\in(1,T)$ as the time steps of Gaussian noise addition. $T$ is the total time step of the forward Markovian diffusion process. $\alpha_{t}(0<\alpha_{t}<1)$ denote the adaptive variables at each iteration and $\gamma_{t}=\prod_{t=0}^{t}\alpha_{t}$. With the above settings, the forward process ($q$: with no learnable parameters) of Mjölnir’s Gradients Diffusion Model can be modeled as:
$q(X_{1}|X_{0})$ $=$ $N(X_{1};\sqrt{\alpha_{1}}X_{0},(1-\alpha_{1})I)$ (8)
$q(X_{t}|X_{0})$ $=$ $N(X_{t};\sqrt{\gamma_{t}}X_{0},(1-\gamma_{t})I)$ (9)
Given $(X_{0},X_{t})$, $X_{t-1}$ can be modeled as:
$q(X_{t-1}|X_{0},X_{t})=N(X_{t};\mu_{t},\sigma_{t}I)$ (10)
If insert the stolen Shared Perturbed Gradients $\nabla W^{\prime}$ or constructed Surrogate Perturbed Gradients $\nabla W^{s}_{perturbed}$ as training condition then input $X_{0}=(condition,\nabla W_{j})$.If not to train with the condition, then input $X_{0}=\nabla W_{j}$.Through algebraic calculation, $\mu_{t}$ and $\sigma_{t}$ in Eq. (10) can be simplified for further usage in the reverse training process(Saharia et al. 2023) as:
$\mu_{t}$ $=$ $\frac{\sqrt{\gamma_{t-1}}(1-\alpha_{t})}{1-\gamma_{t}}X_{0}+\frac{\alpha_{t}(1-\gamma_{t-1})}{1-\gamma_{t}}X_{t}$ (11)
$\sigma^{2}_{t}$ $=$ $\frac{1-\gamma_{t-1}}{1-\gamma_{t}}(1-\alpha_{t})$ (12)
In the reverse process $(p)$ of Mjölnir’s Gradient Diffusion training, the objective function $f_{\theta}$ which is trained to predict noise vector $\epsilon$ is modeled as:
$E_{X_{0},\epsilon}[\frac{(1-\alpha_{t})^{2}}{2\sigma^{2}_{t}\alpha_{t}(1-\gamma_{t})}||f_{\theta}(\sqrt{\gamma_{t}}X_{0}+\sqrt{1-\gamma_{t}}\epsilon,t)-\epsilon||^{2}]$ (13)
Datasets Model $\bm{\varepsilon}$=1 $\bm{\varepsilon}$=5 $\bm{\varepsilon}$=10
$PSNR_{i}$ $LRA$ $PSNR_{i}$ $LRA$ $PSNR_{i}$ $LRA$
MINIST GRNN(Ren, Deng, and Xie 2022) 14.70 1.000 16.54 1.000 20.19 1.000
IG(Geiping et al. 2020) 3.721 - 5.164 - 6.067 -
DLG(Zhu, Liu, and Han 2019) 0.222 0.676 8.025 0.752 17.08 0.909
Mjölnir 17.87 0.851 24.17 0.895 33.09 0.927
Mjölnir(Non-Adaptive) 17.39 0.851 24.14 0.889 32.98 0.925
Mjölnir(Condition) 20.87 0.887 28.39 0.924 35.14 0.957
CIFAR100 GRNN(Ren, Deng, and Xie 2022) 18.24 1.000 20.61 1.000 21.15 1.000
IG(Geiping et al. 2020) 3.684 - 5.073 - 5.971 -
DLG(Zhu, Liu, and Han 2019) 0.112 0.769 7.673 0.833 18.57 0.896
Mjölnir 18.25 0.883 19.04 0.902 21.32 0.902
Mjölnir(Non-Adaptive) 18.15 0.871 17.98 0.884 21.06 0.891
Mjölnir(Condition) 18.69 0.901 21.02 0.902 23.49 0.908
STL10 GRNN(Ren, Deng, and Xie 2022) 12.24 1.000 16.24 1.000 20.98 1.000
IG(Geiping et al. 2020) 4.323 - 6.082 - 7.851 -
DLG(Zhu, Liu, and Han 2019) 0.038 0.783 5.673 0.815 17.25 0.851
Mjölnir 12.36 0.806 16.30 0.857 19.77 0.889
Mjölnir(Non-Adaptive) 12.28 0.801 15.19 0.849 19.76 0.887
Mjölnir(Condition) 12.98 0.816 20.30 0.872 22.77 0.906
Table 1: Privacy leakage capability of Mjölnir variant models and traditional Gradients Leakage Attacks under FL-DP ($\delta$ = $10^{-5}$, $\varepsilon$ = 1, 5, 10). NA: Non-Adaptive; C: Condition; Gray marker: Mjölnir outperforms the highest result of traditional ones.
In Step (4) Recover Original Gradients, refer to Algorithm 3, two different original gradient generation processes are chosen depending on whether or not attackers know the Noise Scale of Perturbation of the Threat Model. Take FL with Gaussian Differential Privacy (Eq. (6)) as an example, consider an experienced and clever attacker who may know the DP Privacy Budget $\varepsilon$ and the probability of information leakage $\delta$ ($\delta$ is likely to be set as $10^{-5}$ from usual practice in DP), combining with the sensitivity $\nabla S$ (which can be estimated if the attacker has some previous information with the target model training dataset), the noise scale can be calculated or estimated to an approximate value $M$ ($M$ defined as the adaptive parameter in [M-Adaptive Process]). So, $\nabla W^{\prime}$ can be modeled as:
$\nabla W^{\prime}=\nabla W+MN(0,I)$ (14)
Considering the forward Markovian gradient diffusion process in Step 3 Eq. (8) &\&& Eq. (9), the relation between stolen perturbed gradients $\nabla W^{\prime}$ and original gradients $\nabla W$ can be remodeled as:
$\frac{1}{\sqrt{1+M^{2}}}\nabla W^{\prime}=\frac{1}{\sqrt{1+M^{2}}}\nabla W+\frac{M}{\sqrt{1+M^{2}}}N(0,I)$ (15)
Recall that our target is to generate $\nabla W^{R}$, so $\frac{1}{\sqrt{1+M^{2}}}\nabla W^{\prime}$ should be considered as $X_{t}$ in the inverse process of Gradient Diffusion Model, while $\nabla W$ stands for $X_{0}$.
Correspondingly, during the construction of the recovered gradient $\nabla W^{R}$, the inverse time steps can be set as $T^{\prime}$. The relationship between $T^{\prime}$ and $M$ is:
$\frac{1}{1+M^{2}}=\prod_{t=0}^{T^{\prime}}\alpha_{t}$ (16)
Since $\gamma_{t}=\prod_{t=0}^{t}\alpha_{t}$ have been predefined and calculated during the forward Markovian diffusion process, $T^{\prime}$ can be fixed in an approximate range $T^{\prime}\in(T^{\prime}_{-},T^{\prime}_{+})$, where $\prod_{t=0}^{T^{\prime}_{-}}\alpha_{t}<\frac{1}{1+M^{2}}<\prod_{t=0}^{T^{\prime}_{+}}\alpha_{t}$ and $T^{\prime}_{-or+}$ are positive integers less than the total forward noise addition time step $T$. On the other hand, if $M$ can not be estimated, which means the attacker knows nothing about the noise scale. Since $\frac{1}{1+M^{2}}\in(0,1)$, the input value of the inference process can simply be set as $c\nabla W^{\prime}$ where $c\in(0,1)$ to adapt the Markovian forward process.
Also, if conditions allow, $c$ can be modeled and predicted by a separate Machine Learning model according to the specific requirement of attackers.
Figure 4: Comparison of the private image recovery procedures to the iterations between Mjölnir variant models and traditional Gradient Leakage Attack methods (DLG(Zhu, Liu, and Han 2019)).
Figure 5: Comparisons on ground truth clients’ private images and corresponding recovered privacy images from Mjölnir variant models and commonly used traditional gradient leakage attacks. ($\delta$ = $10^{-5}$; $\varepsilon=10$)
Experiments
Experimental Setups
(A) Mjölnir Variant Attack Models. Mjölnir (trained with only unperturbed surrogate gradients), Condition Mjölnir (trained with both perturbed gradients and unperturbed surrogate gradients), and Non-Adaptive Mjölnir (without [M-Adaptive Process]: no perturbation scale $M$ as an adaptive parameter during gradient diffusion process) are presented.
(B) Benchmarks and Datasets. We employ MNIST, CIFAR100, and STL10 as client privacy datasets, which also serve as the ground truth for privacy leakage evaluation. We extract the unperturbed original gradients ($\nabla W$) of the aforementioned three datasets from the local training model of the target client as the reference benchmark of gradient denoising under the FL-PP paradigm. The Mjölnir gradient diffusion model is trained with gradients extracted from a separate dataset, FashionMNIST.
(C) Evaluation and Boundary. To evaluate the Privacy Leakage Capability, we utilize the Image Average Peak Signal-to-Noise Ratio $PSNR_{i}$ and the Label Recovered Accuracy $LRA$. These metrics are employed to assess the fidelity of the recovered images and the accuracy of the recovered labels, respectively, to the original images and ground truth labels. The boundary of Privacy Leakage Attack is aligned with previous works: the attack is considered successful if human visual perception can discern the requisite information from the recovered images. In the context of evaluating Gradient Denoising, we employ the cosine similarity $CosSimilar_{g}$ and the Average Peak Signal-to-Noise Ratio $PSNR_{g}$ as metrics to assess the quality of the Recovered Gradients $\nabla W^{R}$ compared to the Original Gradients $\nabla W$. Higher values of the two metrics indicate better accuracy in the original gradient recovery.
Privacy Leakage Capability
The comparison of overall privacy leakage capability from perturbed gradients of Mjölnir and traditional Gradient Leakage Attacks (GRNN (Ren, Deng, and Xie 2022), IG (Geiping et al. 2020), and DLG (Zhu, Liu, and Han 2019)) under FL-DP, compares Image Average Peak Signal Noise Ratio $PSNR_{i}$ and Label Recovered Accuracy $LRA$ of the recovered local clients’ privacy information. Local clients’ private training datasets are MNIST, CIFAR100, and STL10.
According to the numerical experimental results shown in Table 1 and the visualization results shown in Fig. 5, Mjölnir variant models exhibit a substantial superiority over traditional gradient leakage attacks in terms of private image leakage. On average, there is an approximately $209\%$ increase in the recovered image $PSNR_{i}$ when using Mjölnir. For traditional methods, GRNN consistently achieves an LRA metric of 1.000 by using a robust generative network for label recovery, while other methods update images and labels jointly, leading to less accurate results. Moreover, upon examining the private image recovery procedures of Mjölnir variant models and traditional gradient leakage attacks illustrated in Fig. 4, it becomes evident that the attacks incorporating Mjölnir not only achieve considerably improved accuracy in the ultimate reconstruction of private images compared to conventional approaches but also showcase notable advantages in terms of attack iteration rounds and speed (Mjölnir (Condition) $>$ Mjölnir $>$ Mjölnir (Non-Adaptive)). Above all, Mjölnir variant models demonstrate significant advantages in recovering private images.
Model MNIST CIRAF100 STL10
$CosSimilar_{g}$ $PSNR_{g}$ $CosSimilar_{g}$ $PSNR_{g}$ $CosSimilar_{g}$ $PSNR_{g}$
NBNet(Cheng et al. 2021) 0.992 35.74 0.979 27.51 0.979 27.23
SS-BSN(Han and Yu 2023) 0.995 32.35 0.845 22.85 0.845 22.84
AP-BSN(Lee, Son, and Lee 2022) 0.968 29.61 0.892 24.01 0.893 24.12
Mjölnir 0.996 38.76 0.990 30.42 0.990 30.01
Mjölnir (Non-Adaptive) 0.995 38.59 0.990 30.39 0.990 29.87
Mjölnir(Condition) 0.996 38.78 0.990 30.53 0.993 30.22
Table 2: Overall results on gradients denoising under FL-DP. Threat model setting: DP-($\varepsilon$, $\delta$) = (2,$10^{-5}$), with Gaussian perturbation; NA: Non-Adaptive; C: Condition.
Figure 6: Gradients denoising under FL-PP: Gaussian and Laplace perturbed gradients denoising performance of $PSNR_{g}$ and $CosSimilar_{g}$ via different perturbation magnitudes. The same noise magnitude parameter $\varepsilon$ as in FL-DP is used to represent different levels of perturbation. Perturbation decreases, $\varepsilon$ increases. (Private dataset: STL10, $\delta$ = $10^{-5}$)
Gradients Denoising under FL-DP
Among FL-DP, we choose the NbAFL framework (Wei et al. 2020) (Noising before model aggregation FL) as the threat model in this experiment due to its widespread adoption. To ensure a comprehensive evaluation of the effectiveness of the Mjölnir, we train the three Mjölnir variant models, as well as non-diffusion denoising models such as NBNet (Cheng et al. 2021), SS-BSN (Han and Yu 2023), and AP-BSN (Lee, Son, and Lee 2022), using gradients extracted from the surrogate downstream task model trained on the FashionMnist. Following that, we utilize the trained models to denoise the shared perturbed gradients intercepted by the attacker from the target clients. These target clients have locally trained downstream task models using privacy datasets (MNIST, CIFAR100, and STL10). The gradients are protected using the NbAFL(Wei et al. 2020) before being shared with the server.
Based on the experimental results presented in Table 2, it is evident that Mjölnir showcases superior denoising ability for perturbed gradients under FL-DP, with an average Cosine Similarity exceeding 0.992 and a PSNR of 37.68. This represents a significant improvement of over $27\%$ compared to non-diffusion methods. Further analysis reveals that among the three Mjölnir variant models, the performance can be ranked as follows: Mjölnir(Condition)$>$ Mjölnir $>$ Mjölnir (Non-Adaptive). This ranking demonstrates both M-Adaptive Process and conditional training enhance the denoising and generation performance of Mjölnir’s Gradient Diffusion Model. It suggests that selecting an appropriate Mjölnir variant model based on available prior information can improve attack performance.
Model Inference time (s)
NBNet(Cheng et al. 2021) 2.653
SS-BSN(Han and Yu 2023) 26.15
AP-BSN(Lee, Son, and Lee 2022) 7.138
Mjölnir 6.834
Mjölnir (Non-Adaptive) 6.834
Mjölnir(Condition) 6.917
Table 3: Gradient denoising average inference time of Mjölnir variant models and non-diffusion denoising models under FL-PP. (Private dataset: STL10, $\varepsilon$ = 0.5 $\sim$ 10, $\delta$ = $10^{-5}$. Device: NVIDIA GeForce RTX 2060 GPU; Intel(R) Core(TM) i7-10870H CPU at 2.20GHz.)
Gradients Denoising under FL-PP
To evaluate the generalization capability of Mjölnir on gradient denoising, we constructed two different types of noise (Laplace and Gaussian) randomly applied to each layer of the original gradients within the FL-PP framework. The training and inference processes of Mjölnir variant models, and non-diffusion denoising models (NBNet (Cheng et al. 2021), SS-BSN (Han and Yu 2023), and AP-BSN (Lee, Son, and Lee 2022)) are the same as the above experiments on gradients denoising under FL-DP. In contrast to attacks specifically tailored for the FL-DP framework, the gradient denoising experiments on FL-PP showcase Mjölnir’s ability to effectively handle various types of perturbations and adapt to different magnitudes of perturbation.
Referring to the experimental results illustrated in Fig. 6, it is observed that when subjected to Laplace perturbation, Mjölnir variant models exhibit an average improvement of $21.6\%$ in PSNR and $9.2\%$ in Cosine Similarity. Similarly, under Gaussian perturbation, Mjölnir achieves an average enhancement of $21.3\%$ in PSNR and $9.1\%$ in Cosine Similarity. These findings provide compelling evidence for the robustness and stability of the Mjölnir variant models.
We further compared the average inference time of different denoising models under FL-PP. The experimental results presented in Table 3 indicate that Mjölnir variant models exhibit a relative advantage in terms of gradient denoising speed, surpassing non-diffusion methods by an average improvement of $32.7\%$ in inference time.
Limitations: (1) Mjölnir is not effective in attacking perturbations that are not based on gradient diffusion (noise perturbation) such as representation perturbation. (2) The overall effectiveness of Mjölnir in reconstructing privacy datasets is also bounded by the selected subsequent Gradient Leakage Attacks. (3) Mjölnir is used to attack CNN and (or) DNN-based models. Experimental results suggest that it struggles with transformer-based models due to their complex structure and larger number of parameters, making it difficult to accurately denoise gradients and recover data.
Defense Strategies:Possible defense strategies against Mjölnir can be approached by preserving the privacy of original data, preserving the target attack models, and shared gradients protection. Mjölnir effectively leakages privacy data by attacking the perturbed shared gradients, which suggests the possible defense approaches against Mjölnir should either be non-perturbation gradient protection methods or non-gradient privacy-preserving methods. For instance, processing of the participant’s local training model in depth and layers to weaken the performance of Gradient Leakage Attacks, encryption on original local privacy data before training, and various cryptography-based protections on shared gradients (Huang et al. 2021; Zhang et al. 2022; Rodríguez-Barroso et al. 2023). Cryptography-based approaches always cause more computation time and storage (Zhang et al. 2020), thus lightweight defense strategies that can serve as substitutes for invalid perturbation protection still need future explorations.
Conclusion
This paper makes the first attempt to investigate the diffusion property of the widely used perturbation-based gradient protection. To reveal potential vulnerabilities, we propose a novel Perturbation-Resilient Gradient Leakage Attack via an adaptive diffusion process. This effective attack paradigm deactivates perturbation protection by leveraging the denoising capability of diffusion models without access to clients’ models and external data. Based on Mjölnir, we wish to enhance public consciousness of the privacy leakage issues of existing perturbation-based defenses on gradients.
Checklist
This paper:
  • Includes a conceptual outline and/or pseudocode description of AI methods introduced (yes/partial/no/NA) yes
  • Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results (yes/no) yes
  • Provides well marked pedagogical references for less-familiare readers to gain background necessary to replicate the paper (yes/no) yes
Does this paper make theoretical contributions? (yes/no). If yes, please complete the list below yes
  • All assumptions and restrictions are stated clearly and formally. (yes/partial/no) yes
  • All novel claims are stated formally (e.g., in theorem statements). (yes/partial/no) yes
  • Proofs of all novel claims are included. (yes/partial/no) yes
  • Proof sketches or intuitions are given for complex and/or novel results. (yes/partial/no) yes
  • Appropriate citations to theoretical tools used are given. (yes/partial/no) yes
  • All theoretical claims are demonstrated empirically to hold. (yes/partial/no/NA) yes
  • All experimental code used to eliminate or disprove claims is included. (yes/no/NA) yes
Does this paper rely on one or more datasets? (yes/no). If yes, please complete the list below. yes
  • A motivation is given for why the experiments are conducted on the selected datasets (yes/partial/no/NA) yes
  • All novel datasets introduced in this paper are included in a data appendix. (yes/partial/no/NA) NA
  • All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. (yes/partial/no/NA) NA
  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations. (yes/no/NA) yes
  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available. (yes/partial/no/NA) yes
  • All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing. (yes/partial/no/NA) NA
Does this paper include computational experiments? (yes/no). If yes, please complete the list below. yes
  • Any code required for pre-processing data is included in the appendix. (yes/partial/no). yes
  • All source code required for conducting and analyzing the experiments is included in a code appendix. (yes/partial/no) yes
  • All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. (yes/partial/no) yes
  • All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from (yes/partial/no) yes
  • If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results. (yes/partial/no/NA) yes
  • This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks. (yes/partial/no) yes
  • This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics. (yes/partial/no) yes
  • This paper states the number of algorithm runs used to compute each reported result. (yes/no) yes
  • Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information. (yes/no) yes
  • The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank). (yes/partial/no) yes
  • This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments. (yes/partial/no/NA) yes
  • This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting. (yes/partial/no/NA) yes